You have a 500-foot roll of fencing and a large field. You want to construct a rectangular playground area. What are the dimensions of the largest such yard? What is the largest area?
Answer by edjones(8007)

A square is the largest such yard.
500/4=125' by 125' dimensions of the largest such yard.
Area=125^2=15,625 ft^2
.
Ed